The sum of the first n squares, given by the sum of the terms 1+4+9+...n^2 can be calculated by using the formula n/6(n+1)(2n+1).
What is the value of the 20th term of the series?
Find the sum of the first 25 squares.
Do I just have to put 20 and 25 instead of n? Cos I'm not quite sure what I have to do.

Answer by richwmiller(17219)

to find the 20th term
a20=20^2=400
The sum of the first 25 terms is the formula given
n/6(n+1)(2n+1) where n =25
25/6(25+1)(2*25+1)
25/6*26*51
25/(2*3)*2*13*3*17
25*13*17=5525